OpenStudy (anonymous):

simplify, if the quotient is undefined state this. show work. (-2/7) (5/-8) is the answer 5/28.

hero (hero):

-10/-56 = 10/56 = 5/28 You are correct.

OpenStudy (anonymous):

thanks